OpenStudy (anonymous):

what is th evalue of x in the proportion x over x+10 = 7over 12

OpenStudy (anonymous):

\[\frac{x}{x+10}=\frac{7}{12}\] invert \[\frac{x+10}{x}=\frac{12}{7}\] \[1+\frac{10}{x}=12/7\]

OpenStudy (hoblos):

cross multiply... 12x = 7x +70 5x=70 x=14